The D string of a guitar is stretched to a tension of 150 N. The string has a mass of 3.4 g and is stretched to a length of 0.69 m. What is the speed of waves on the string


1
Expert's answer
2022-01-11T04:13:56-0500

The linear density of the string is the following:


"\mu = \dfrac{m}{L}"

where "m = 3.4g = 3.4\times 10^{-3}kg" and "L = 0.69m". The speed of waves is given then as follows:


"v = \sqrt{\dfrac{T}{\mu}} = \sqrt{\dfrac{TL}{m}}"

where "T = 150N".

Thus, obtain:


"v = \sqrt{\dfrac{150N\cdot 0.69m}{3.4\times 10^{-3}kg}} \approx 1.7\times 10^2m/s"

Answer. "1.7\times 10^2m/s".